# gRNAde: Geometric Deep Learning for 3D RNA inverse design

### v4

@@ 1 Introduction
Challenges of RNA modelling. Despite the promises of RNA therapeutics, proteins have instead been the primary focus in the 3D biomolecular modelling community. Availability of large-scale protein structures from the PDB combined with advances in deep learning for structured data (Bronstein et al., 2021; Duval et al., 2023) have revolutionized protein 3D structure prediction (Jumper et al., 2021) and rational design (Dauparas et al., 2022; Watson et al., 2023). Applications of deep learning for computational RNA design are underexplored compared to proteins due to paucity of 3D structural data (Schneider et al., 2023). Most tools for RNA design primarily focus on secondary structure without considering 3D geometry (Churkin et al., 2018) and use non-learnt algorithms for aligning 3D RNA fragments (Han et al., 2017; Yesselman et al., 2019), which can be restrictive due to the hand-crafted nature of the heuristics used.
@@
  • Enables multi-state and partial RNA design, which were previously not possible with Rosetta. gRNAde with multi-state GNNs moderately improves sequence recovery over an equivalent single-state model on a new benchmark of structurally flexible RNAs, especially for surface nucleotides which undergo positional or secondary structural changes.
@@ 2.2 RNA conformational ensembles as geometric multi-graphs
Featurization. The input to gRNAde is a set of 3D RNA backbone structures (a conformational ensemble) and the corresponding sequence of $n$ nucleotides, for instance via PDB files. As shown in Figure 2, gRNAde builds a geometric graph representation for each input structure:
@@
    We start with a 3-bead coarse-grained representation of the RNA backbone, retaining the coordinates for P, C4’, N1 (pyrimidine) or N9 (purine) for each nucleotide (Dawson et al., 2016). This ‘pseudotorsional’ representation can describe RNA backbone conformations completely in most cases while reducing the size of the torsional space (Wadley et al., 2007).
@@
    Nodes are initialized with geometric features analogous to the featurization used in protein inverse folding (Ingraham et al., 2019; Jing et al., 2020): (a) forward and reverse unit vectors along the backbone backbone from the 5’ end to the 3’ end, ($\vec{{\bm{x}}}_{i+1}-\vec{{\bm{x}}}_{i}$ and $\vec{{\bm{x}}}_{i}-\vec{{\bm{x}}}_{i-1}$); and (b) unit vectors, distances, angles, and torsions from each C4’ to the corresponding P and N1/N9.
@@
Multi-graph representation. As described in the previous section, given a set of $k$ structures in the input conformational ensemble, each RNA backbone is featurized as a separate geometric graph $\mathcal{G}^{(k)}=({\bm{A}}^{(k)},{\bm{S}}^{(k)},\vec{{\bm{V}}}^{(k)})$ with the scalar features ${\bm{S}}^{(k)}\in\mathbb{R}^{n\times f}$, vector features $\vec{{\bm{V}}}^{(k)}\in\mathbb{R}^{n\times f^{\prime}\times 3}$, and ${\bm{A}}^{(k)}$, an $n\times n$ adjacency matrix. For clear presentation and without loss of generality, we omit edge features and use $f$, $f^{\prime}$ to denote scalar/vector feature channels.
@@ 2.3 Multi-state GNN for representation learning on conformational ensembles
The gRNAde model, illustrated in Figure 3, processes one or more RNA backbone graphs (a conformational ensemble) via a multi-state GNN encoder which is equivariant to 3D roto-translation of coordinates as well as to the ordering of conformers, followed by conformer order-invariant pooling and sequence decoding. We describe each component in the following sections.
Multi-state GNN encoder. When representing conformational ensembles as a multi-graph, each node feature tensor contains three axes: (no. of nodes, no. of conformations, feature channels). We perform message passing on the multi-graph adjacency to independently process each conformer, while maintaining permutation equivariance of the updated feature tensors along both the first (no. of nodes) and second (no. of conformations) axes. This works by operating on only the feature channels axis and generalising the PyTorch Geometric (Fey and Lenssen, 2019) message passing class to account for the extra conformations axis; see Figure 10 and the pseudocode in the Appendix.
@@
A simple sum or average pooling does not introduce any new learnable parameters to the pipeline and is flexible to handle a variable number of conformations, enabling both single-state and multi-state design. It may be interesting to ablate the impact of more expressive geometric GNN layers (Joshi et al., 2023) as well as set pooling functions (Maron et al., 2020) in future iterations of gRNAde.
@@
Sampling. When using gRNAde for inference and designing new sequences, we iteratively sample the base identity for a given nucleotide from the predicted conditional probability distribution, given the partially designed sequence up till that nucleotide/decoding step. We can modulate the smoothness or sharpness of the probability distribution by using a temperature parameter. At lower temperatures, for instance $\leq$1.0, we expect higher native sequence recovery and lower diversity in gRNAde’s designs. At higher temperatures, the model produces more diverse designs by sampling from a smoothed probability distribution. We can also consider unordered decoding (Dauparas et al., 2022) and masking or logit biasing during sampling, depending on the design scenario at hand. This enables gRNAde to perform partial re-design of RNA sequences, retaining specified nucleotide identities while designing the rest of the sequence. Similar approaches for functional protein design have been shown to be successful in the wet lab (Sumida et al., 2024).
@@ 2.4 Evaluation metrics for designed sequences
In principle, inverse folding models can be sampled from to obtain a large number of designed sequences for a given backbone structure. Thus, in-silico metrics to determine which sequences are useful and which ones to prioritise in wet lab experiments are a critical part of the overall pipeline. We currently use the following metrics to evaluate gRNAde’s designs, visualised in Appendix Figure 8:
  • Native sequence recovery, which is the average percentage of native (ground truth) nucleotides correctly recovered in the sampled sequences. Recovery is the most widely used metric for biomolecule inverse design (Dauparas et al., 2022).
  • Secondary structure self-consistency score, where we ‘forward fold’ the sampled sequences using a secondary structure prediction tool (we used EternaFold (Wayment-Steele et al., 2022)) and measure the average Mathew’s Correlation Coefficient (MCC) to the ground truth secondary structure, represented as a binary adjacency matrix. MCC values range between -1 and +1, with +1 represents a perfect prediction, 0 an average random prediction and -1 an inverse prediction. Similar forward folding metrics, sometimes termed ‘designability’, have been found to correlate with experimental success in protein design (Watson et al., 2023)111 It would be ideal to consider a 3D structure self-consistency score, instead. While RNA 3D structure prediction is currently not at the same level of reliability as that for proteins, our framework is flexible enough to include new self-consistency scores based on improved RNA structure prediction tools in the future. .
@@ 3 Experimental Setup
3D RNA structure dataset. We create a machine learning-ready dataset for RNA inverse design using RNASolo (Adamczyk et al., 2022), a novel repository of RNA 3D structures extracted from solo RNAs, protein-RNA complexes, and DNA-RNA hybrids in the PDB. We used structures at resolution $\leq$4.0Å  resulting in 4,223 unique RNA sequences for which a total of 12,011 structures are available (downloaded from RNASolo on 31 October 2023). Dataset statistics are available in Figure 9 in the Appendix, illustrating the diversity of our dataset in terms of sequence length, number of structures per sequence, as well as structural variations among conformations per sequence.
Structural clustering. In order to ensure that we evaluate gRNAde’s generalization ability to novel RNAs, we cluster the 4,223 unique RNAs into groups based on structural similarity. We use qTMclust for efficiently applying US-align (Zhang et al., 2022) with a similarity threshold of TM-score > 0.45, and ensure that we train, validate and test gRNAde on structurally dissimilar clusters (see next paragraph). We also provide utilities for clustering based on sequence homology using CD-HIT (Fu et al., 2012), which leads to splits containing biologically dissimilar clusters of RNAs.
@@
    Single-state split. This split is used to fairly evaluate gRNAde for single-state design on a set of RNA structures of interest from the PDB identified by Das et al. (2010), which mainly includes riboswitches, aptamers, and ribozymes (full listing in Table 1). We identify the structural clusters belonging to the RNAs identified in Das et al. (2010) and add all the RNAs in these clusters to the test set (100 samples). The remaining clusters are randomly added to the training and validation splits.
@@
    Multi-state split. This split is used to tests gRNAde’s ability to design RNA with multiple distinct conformational states. We order the structural clusters based on median intra-sequence RMSD among available structures within the cluster [^2] For each RNA sequence, we compute the pairwise C4’ RMSD among all available structures. We then compute the median RMSD across all sequences within each structural cluster.. The top 100 samples from clusters with the highest median intra-sequence RMSD are added to the test set. The next 100 samples are added to the validation set and all remaining samples are used for training.
@@
Evaluation metrics. For a given data split, we evaluate models on the held-out test set by designing 16 sequences (sampled at temperature 0.1) for each test data point and computing averages for each of the three metrics described in Section 2.4: native sequence recovery, secondary structure MCC self-consistency score, and perplexity. We employ early stopping by reporting test set performance for the model checkpoint for the epoch with the best validation set recovery. Standard deviations are reported across 3 consistent random seeds for all models.
Hyperparameters. All models use 4 encoder and 4 decoder GVP-GNN layers, with 128 scalar/16 vector node features, 64 scalar/4 vector edge features, and drop out probability 0.5, resulting in 2,147,944 trainable parameters. All models are trained for a maximum of 50 epochs using the Adam optimiser with an initial learning rate of 0.0001, which is reduced by a factor 0.9 when validation performance plateaus with patience of 5 epochs. Ablation studies of key modelling decisions are available in Appendix Table 2.
4 Results
@@
Figure 4: gRNAde compared to Rosetta for single-state design. (a) We benchmark native sequence recovery of gRNAde, Rosetta, and FARNA on 14 RNA structures of interest identified by Das et al. (2010). gRNAde obtains higher native sequence recovery rates (56% on average) compared to Rosetta (45%) and FARNA (32%). (b) Sequence recovery per sample for Rosetta and gRNAde, shaded by gRNAde’s perplexity for each sample. gRNAde’s perplexity is correlated with native sequence recovery for designed sequences. Full results available in Appendix Table 1.
@@ 4.1 Single-state RNA design benchmark
gRNAde improves sequence recovery over Rosetta. In Figure 4, we compare gRNAde’s native sequence recovery for single-state design with numbers taken from Das et al. (2010) for Rosetta and FARNA (a predecessor of Rosetta). gRNAde has higher recovery of 56% on average compared to 45% for Rosetta and 32% for FARNA. See Appendix Table 1 for full results.
gRNAde is significantly faster than Rosetta. In addition to superior sequence recovery, gRNAde is significantly faster than Rosetta for high-throughout design pipelines. Training gRNAde from scratch takes roughly 2–6 hours on a single A100 GPU, depending on the exact hyperparameters. Once trained, gRNAde can design hundreds of sequences for backbones with hundreds of nucleotides in $\sim$1 second with GPU acceleration. On the other hand, Rosetta takes order of hours to produce a single design due performing expensive Monte Carlo optimisations333 While we have not run Rosetta ourselves, we note that its documentation states that “runs on RNA backbones longer than  ten nucleotides take many minutes or hours”. . Deep learning methods like gRNAde are arguably easier to use since no expert customization is required and setup is easier compared to Rosetta (Dauparas et al., 2022), potentially making RNA design more broadly accessible.
gRNAde’s perplexity correlates with recovery. In Figure 4(b), we plot native sequence recovery per sample for Rosetta vs. gRNAde, shaded by gRNAde’s average perplexity for each sample. Perplexity is an indicator of the model’s confidence in its own prediction (lower perplexity implies higher confidence) and appears to be correlated with native sequence recovery. In Section 4.3, we further demonstrate the utility of gRNAde’s perplexity for unsupervised learning of RNA fitness landscapes.
Figure 5: Multi-state design benchmark. Multi-state gRNAde show marginal improvement over an equivalent single-state model in terms of average native sequence recovery and secondary structure self-consistency score. We plot performance for one consistent random seed across all models; collated results are available in Appendix Table 2.
@@ 4.2 Multi-state RNA design benchmark
Structured RNAs often adopt multiple distinct conformational states to perform biological functions (Ken et al., 2023). For instance, riboswitches adopt at least two distinct functional conformation: a ligand bound (holo) and unbound (apo) state, which helps them regulate and control gene expression (Stagno et al., 2017). If we were to attempt single-state inverse design for such RNAs, each backbone structure may lead to a different set of sampled sequences. It is not obvious how to select the input backbone as well as designed sequence when using single-state models for multi-state design444 ProteinMPNN (Dauparas et al., 2022) proposes to average logits from multiple backbones for multi-state protein design. Here is a simple example to highlight issues with such an approach: Consider two states A and B, and choice of labels X, Y, and Z. For state A: X, Y, Z are assigned probabilities 75%, 20%, 5%. For state B: X, Y, Z are assigned probabilities 5%, 20%, 75%. Logically, label Y is the only one that is compatible with both states. However, averaging the probabilities would lead to label X or Z being more likely to be sampled in designs. . gRNAde’s multi-state GNN, descibed in Section 2.3, directly ‘bakes in’ the multi-state nature of RNA as an architectural inductive bias and designs sequences explicitly conditioned on multiple states.
@@
Multi-state gRNAde boosts sequence recovery. In Figure 5, we compared a single-state variant of gRNAde with otherwise equivalent multi-state models (with up to 3 and 5 states, respectively) in terms of native sequence recovery and secondary structure self-consistency score555 Self-consistency scores for multi-state design are perhaps less reliable than single-state design due to the lack of multi-state structure prediction tools. At present, we compute the average MCC score between the multiple ground truth secondary structures and the predicted structure for designed sequences. . Multi-state variants show marginal improvements for both metrics. As a caveat, it is worth noting that multi-state models consume more GPU memory than an equivalent single-state model during mini-batch training (approximate peak GPU usage for max. number of states set to 1: 12GB, 3: 28GB, 5: 50GB on a single A100 with otherwise equivalent hyperparameters).
Figure 6: Per-nucleotide sequence recovery for multi-state design. Multi-state gRNAde shows improved sequence recovery over a single-state model for structurally flexible regions of RNAs, as characterised by nucleotides that tend to undergo changes in base pairing (left), nucleotides with greater average solvent accessible surface area (centre), and nucleotides with higher average RMSD (right) across multiple states. Marginal histograms in blue show the distribution of values.
Improved recovery in structurally flexible regions. In Figure 6, we evaluated gRNAde’s multi-state sequence recovery at a fine-grained, per-nucleotide level. gRNAde with multi-state GNN architectures improves sequence recovery over its single-state variant on structurally flexible nucleotides, as characterised by undergoing changes in base pairing/secondary structure, higher average RMSD between 3D coordinates across states, and larger solvent accessible surface area.

### v5

@@ 1 Introduction
Challenges of RNA modelling. Despite the promises of RNA therapeutics, proteins have instead been the primary focus in the 3D biomolecular modelling community. Availability of a large number of protein structures from the PDB combined with advances in deep learning for structured data (Bronstein et al., 2021; Duval et al., 2023) have revolutionized protein 3D structure prediction (Jumper et al., 2021) and rational design (Dauparas et al., 2022; Watson et al., 2023). Applications of deep learning for computational RNA design are underexplored compared to proteins due to paucity of 3D structural data (Schneider et al., 2023). Most tools for RNA design primarily focus on secondary structure without considering 3D geometry (Churkin et al., 2018) and use non-learnt algorithms for aligning 3D RNA fragments (Han et al., 2017; Yesselman et al., 2019), which can be restrictive due to the hand-crafted nature of the heuristics used.
@@
  • Enables multi-state RNA design, which was previously not possible with Rosetta. gRNAde with multi-state GNNs improves sequence recovery over an equivalent single-state model on a benchmark of structurally flexible RNAs, especially for surface nucleotides which undergo positional or secondary structural changes.
@@ 2.2 RNA conformational ensembles as geometric multi-graphs
Featurization. The input to gRNAde is an RNA to be re-designed. For instance, this could be a set of PDB files with 3D backbone structures for the given RNA (a conformational ensemble) and the corresponding sequence of $n$ nucleotides. As shown in Figure 2, gRNAde builds a geometric graph representation for each input structure:
@@
    We start with a 3-bead coarse-grained representation of the RNA backbone, retaining the coordinates for P, C4’, N1 (pyrimidine) or N9 (purine) for each nucleotide (Dawson et al., 2016). This ‘pseudotorsional’ representation describes RNA backbones completely in most cases while reducing the size of the torsional space to prevent overfitting (Wadley et al., 2007).
@@
    Nodes are initialized with geometric features analogous to the featurization used in protein inverse folding (Ingraham et al., 2019; Jing et al., 2020): (a) forward and reverse unit vectors along the backbone from the 5’ end to the 3’ end, ($\vec{{\bm{x}}}_{i+1}-\vec{{\bm{x}}}_{i}$ and $\vec{{\bm{x}}}_{i}-\vec{{\bm{x}}}_{i-1}$); and (b) unit vectors, distances, angles, and torsions from each C4’ to the corresponding P and N1/N9.
@@
Multi-graph representation. As described in the previous section, given a set of $k$ (conformer) structures in the input conformational ensemble, each RNA backbone is featurized as a separate geometric graph $\mathcal{G}^{(k)}=({\bm{A}}^{(k)},{\bm{S}}^{(k)},\vec{{\bm{V}}}^{(k)})$ with the scalar features ${\bm{S}}^{(k)}\in\mathbb{R}^{n\times f}$, vector features $\vec{{\bm{V}}}^{(k)}\in\mathbb{R}^{n\times f^{\prime}\times 3}$, and ${\bm{A}}^{(k)}$, an $n\times n$ adjacency matrix. For clear presentation and without loss of generality, we omit edge features and use $f$, $f^{\prime}$ to denote scalar/vector feature channels.
@@ 2.3 Multi-state GNN for representation learning on conformational ensembles
The gRNAde model, illustrated in Appendix Figure 13, processes one or more RNA backbone graphs via a multi-state GNN encoder which is equivariant to 3D roto-translation of coordinates as well as to the ordering of conformers, followed by conformer order-invariant pooling and sequence decoding. We describe each component in the following sections.
Multi-state GNN encoder. When representing conformational ensembles as a multi-graph, each node feature tensor contains three axes: (#nodes, #conformations, feature channels). We perform message passing on the multi-graph adjacency to independently process each conformer, while maintaining permutation equivariance of the updated feature tensors along both the first (#nodes) and second (#conformations) axes. This works by operating on only the feature channels axis and generalising the PyTorch Geometric (Fey and Lenssen, 2019) message passing class to account for the extra conformations axis; see Appendix Figure 14 and the pseudocode for details.
@@
A simple sum or average pooling does not introduce any new learnable parameters to the pipeline and is flexible to handle a variable number of conformations, enabling both single-state and multi-state design with the same model.
@@
Sampling. When using gRNAde for inference and designing new sequences, we iteratively sample the base identity for a given nucleotide from the predicted conditional probability distribution, given the partially designed sequence up until that nucleotide/decoding step. We can modulate the smoothness or sharpness of the probability distribution by using a temperature parameter. At lower temperatures, for instance $\leq$1.0, we expect higher native sequence recovery and lower diversity in gRNAde’s designs. At higher temperatures, the model produces more diverse designs by sampling from a smoothed probability distribution. We can also consider unordered decoding (Dauparas et al., 2022) and masking or logit biasing during sampling, depending on the design scenario at hand. This enables gRNAde to perform partial re-design of RNA sequences, retaining specified nucleotide identities while designing the rest of the sequence. Similar approaches for functional protein design have been shown to be successful in the wet lab (Sumida et al., 2024).
Figure 3: In-silico evaluation metrics for gRNAde designed sequences. We consider (1) sequence recovery, the percentage of native nucleotides recovered in designed samples, (2) self-consistency scores, which are measured by ‘forward folding’ designed sequences using a structure predictor and measuring how well 2D and 3D structure are recovered (we use EternaFold and RhoFold for 2D/3D structure prediction, respectively). We also report (3) perplexity, the model’s estimate of the likelihood of a sequence given a backbone.
@@ 2.4 Evaluation metrics for designed sequences
In principle, inverse folding models can be sampled from to obtain a large number of designed sequences for a given backbone structure. Thus, in-silico metrics to determine which sequences are useful and which ones to prioritise in wet lab experiments are a critical part of the overall pipeline. We currently use the following metrics to evaluate gRNAde’s designs, visualised in Figure 3:
  • Native sequence recovery, which is the average percentage of native (ground truth) nucleotides correctly recovered in the sampled sequences. Recovery is the most widely used metric for biomolecule inverse design (Dauparas et al., 2022) but can be misleading in the case of RNAs where alternative nucleotide base pairings can form the same structural patterns.
  • Secondary structure self-consistency score, where we ‘forward fold’ the sampled sequences using a secondary structure prediction tool (we used EternaFold (Wayment-Steele et al., 2022)) and measure the average Matthew’s Correlation Coefficient (MCC) to the groundtruth secondary structure, represented as a binary adjacency matrix. MCC values range between -1 and +1, where +1 represents a perfect prediction, 0 an average random prediction and -1 an inverse prediction. This measures how well the designs recover base pairing patterns.
  • Tertiary structure self-consistency scores, where we ‘forward fold’ the sampled sequences using a 3D structure prediction tool (we used RhoFold (Shen et al., 2022)) and compute the average RMSD, TM-score and GDT_TS to the groundtruth C4’ coordinates to measure how well the designs recover global structural similarity and 3D conformations.
@@ 3 Experimental Setup
3D RNA structure dataset. We create a machine learning-ready dataset for RNA inverse design using RNASolo (Adamczyk et al., 2022), a novel repository of RNA 3D structures extracted from solo RNAs, protein-RNA complexes, and DNA-RNA hybrids in the PDB. We used structures at resolution $\leq$4.0Å  resulting in 4,223 unique RNA sequences for which a total of 12,011 structures are available (RNASolo date cutoff: 31 October 2023). Dataset statistics are available in Appendix Figure 15, illustrating the diversity of our dataset in terms of sequence length, number of structures per sequence, as well as structural variations among conformations per sequence.
Structural clustering. In order to ensure that we evaluate gRNAde’s generalization ability to novel RNAs, we cluster the 4,223 unique RNAs into groups based on structural similarity. We use US-align (Zhang et al., 2022) with a similarity threshold of TM-score >0.45 for clustering, and ensure that we train, validate and test gRNAde on structurally dissimilar clusters (see next paragraph). We also provide utilities for clustering based on sequence homology using CD-HIT (Fu et al., 2012), which leads to splits containing biologically dissimilar clusters of RNAs.
@@
    Single-state split. This split is used to fairly evaluate gRNAde for single-state design on a set of RNA structures of interest from the PDB identified by Das et al. (2010), which mainly includes riboswitches, aptamers, and ribozymes. We identify the structural clusters belonging to the RNAs identified in Das et al. (2010) and add all the RNAs in these clusters to the test set (100 samples). The remaining clusters are randomly added to the training and validation splits.
@@
    Multi-state split. This split is used to test gRNAde’s ability to design RNA with multiple distinct conformational states. We order the structural clusters based on median intra-sequence RMSD among available structures within the cluster [^1] For each RNA sequence, we compute the pairwise C4’ RMSD among all available structures. We then compute the median RMSD across all sequences within each structural cluster.. The top 100 samples from clusters with the highest median intra-sequence RMSD are added to the test set. The next 100 samples are added to the validation set and all remaining samples are used for training.
@@
Evaluation metrics. For a given data split, we evaluate models on the held-out test set by designing 16 sequences (sampled at temperature 0.1) for each test data point and computing averages for each of the metrics described in Section 2.4: native sequence recovery, structural self-consistency scores and perplexity. We employ early stopping by reporting test set performance for the model checkpoint for the epoch with the best validation set recovery. Standard deviations are reported across 3 consistent random seeds for all models.
Hyperparameters. All models use 4 encoder and 4 decoder GVP-GNN layers, with 128 scalar/16 vector node features, 64 scalar/4 vector edge features, and drop out probability 0.5, resulting in 2,147,944 trainable parameters. All models are trained for a maximum of 50 epochs using the Adam optimiser with an initial learning rate of 0.0001, which is reduced by a factor 0.9 when validation performance plateaus with patience of 5 epochs. Detailed ablation studies of key modelling decisions are available in Appendix D.
@@
Figure 4: gRNAde compared to Rosetta for single-state design. (a) We benchmark native sequence recovery of gRNAde, Rosetta, FARNA and ViennaRNA on 14 RNA structures of interest identified by Das et al. (2010). gRNAde obtains higher native sequence recovery rates (56% on average) compared to Rosetta (45%). (b) Sequence recovery per sample for Rosetta and gRNAde, shaded by gRNAde’s perplexity for each sample. gRNAde’s perplexity is correlated with native sequence recovery for designed sequences. Full results are available in Appendix Table 2.
4 Results
@@ 4.1 Single-state RNA design benchmark
gRNAde improves sequence recovery over Rosetta. In Figure 4, we compare gRNAde’s native sequence recovery for single-state design with numbers taken from Das et al. (2010) for Rosetta, FARNA (a predecessor of Rosetta), and ViennaRNA (the most popular 2D inverse folding method). gRNAde has higher recovery of 56% on average compared to 45% for Rosetta, 32% for FARNA, and 27% for ViennaRNA. See Appendix Table 2 for per-RNA results.
[…]
gRNAde’s perplexity correlates with sequence and structural recovery. In Figure 4(b), we plot native sequence recovery per sample for Rosetta vs. gRNAde, shaded by gRNAde’s average perplexity for each sample. Perplexity is an indicator of the model’s confidence in its own prediction (lower perplexity implies higher confidence) and appears to be correlated with native sequence recovery. Additionally, visualisations of gRNAde’s designs for a riboswitch in Figure 5 show that perplexity is also correlated with structural self-consistency scores. In the subsequent Section 4.3, we further demonstrate the utility of gRNAde’s perplexity for zero-shot ranking of RNA fitness landscapes.
Design 1:
perplexity: 1.310
recovery: 0.591 (27 edits)
sc2D = 0.923, scRMSD = 1.384
scTM = 0.831, scGDT = 0.830
Design 2:
perplexity: 1.382
recovery: 0.409 (37 edits)
sc2D = 0.922, scRMSD = 2.125
scTM = 0.687, scGDT = 0.678
Design 3:
perplexity: 1.425
recovery: 0.515 (30 edits)
sc2D = 0.923, scRMSD = 3.213
scTM = 0.512, scGDT = 0.526
Figure 5: Cherry-picked designs for Guanine riboswitch aptamer (PDB: 4FE5). We show the RhoFold-predicted 3D structure in colour overlaid on the groundtruth structure in grey. Designs recover the base pairing patterns and tertiary structure of the RNA, as measured by high self-consistency score. gRNAde’s perplexity is correlated well with 3D self-consistency scores and can be useful for ranking designs. More design visualisations are available in Appendix A.
@@ 4.2 Multi-state RNA design benchmark
Structured RNAs often adopt multiple distinct conformational states to perform biological functions (Ken et al., 2023). For instance, riboswitches adopt at least two distinct functional conformations: a ligand bound (holo) and unbound (apo) state, which helps them regulate and control gene expression (Stagno et al., 2017). If we were to attempt single-state inverse design for such RNAs, each backbone structure may lead to a different set of sampled sequences. It is not obvious how to select the input backbone as well as designed sequence when using single-state models for multi-state design. gRNAde’s multi-state GNN, descibed in Section 2.3, directly ‘bakes in’ the multi-state nature of RNA into the architecture and designs sequences explicitly conditioned on multiple states.
@@
Multi-state gRNAde boosts sequence recovery. In Figure 6(a), we compared a single-state variant of gRNAde with otherwise equivalent multi-state models (with up to 3 and 5 states, respectively) in terms of native sequence recovery  [^3] Self-consistency scores for multi-state design are perhaps less reliable than single-state design due to the lack of multi-state structure prediction tools. At present, we compute the average score between multiple ground truth structures and the predicted structure for designed sequences.. Multi-state variants show marginal improvements, overall. As a caveat, it is worth noting that multi-state models consume more GPU memory than an equivalent single-state model during mini-batch training (approximate peak GPU usage for max. number of states = 1: 12GB, 3: 28GB, 5: 50GB on a single A100 with at most 3000 total nodes in a mini-batch).
Improved recovery in structurally flexible regions. In Figure 6(b), we evaluated gRNAde’s multi-state sequence recovery at a fine-grained, per-nucleotide level. Multi-state GNNs improve sequence recovery over the single-state variant on structurally flexible nucleotides, as characterised by undergoing changes in base pairing/secondary structure, higher average RMSD between 3D coordinates across states, and larger solvent accessible surface area.
(a) Per-sample sequence recovery
(b) Per-nucleotide recovery vs. structural flexibility measures
Figure 6: Multi-state design benchmark. (a) Multi-state gRNAde show marginal improvement over an equivalent single-state model in terms of average per-sample sequence recovery over all test RNAs. (b) When plotting sequence recovery per-nucleotide, multi-state gRNAde improves over a single-state model for structurally flexible regions of RNAs, as characterised by nucleotides that tend to undergo changes in base pairing (left), nucleotides with greater average solvent accessible surface area (centre), and nucleotides with higher average RMSD (right) across multiple states. Marginal histograms in blue show the distribution of values. We plot performance for one consistent random seed across all models; collated results and ablations are available in Appendix Table 1.